OpenStudy (animelover91804):

What value of j makes the statement true? j – 13.7 = –6.5

OpenStudy (animelover91804):

XDDD

OpenStudy (alexandervonhumboldt2):

add 13.7 to both sides

OpenStudy (alexandervonhumboldt2):

j-13.7+13.7=-6.5+13.7 j=-6.5+13.7=???